OpenStudy (anonymous):

Solve for x: 3/3x+1/x+410/7x

OpenStudy (anonymous):

i got x= 1/4 andx=-2 is that right

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@Paynesdad

hero (hero):

\[\frac{3}{3x + 1} \div \frac{x + 410}{7x}\] Reciprocate the second fraction to write the expression as the multiplication of two fractions, then simplify.

OpenStudy (anonymous):

so i take x out of the numerator and denominator

OpenStudy (anonymous):

i an so lost on this question

OpenStudy (anonymous):

@helpplzzzz12120 You can't solve for x if there is no equal sign!

hero (hero):

Good observation @Paynesdad

OpenStudy (anonymous):

oh i stink with fractions like this

OpenStudy (anonymous):

Fractions is tough for lots of students. Too many students weren't really taught well and rigorously when they should have been in like 3rd -5th grade ("because they have calculators")...and now they suffer in higher level math. I made my son learn fractions backwards and forwards when he was young so he wouldn't struggle so much when he got to higher level math.

OpenStudy (anonymous):

@helpplzzzz12120 Still the fact remains with out an equals sign there is no way to solve for x ... check the question and see if you posted on here properly.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5/(x^2-1-2(x)=2/(x+1)

OpenStudy (anonymous):

That looks totally different!

OpenStudy (anonymous):

\[5divx^2-1-2divx=2divx+1\]

OpenStudy (anonymous):

I am not sure what that equation means ... bear with me. Is that \[\frac{5}{x^2-1}-2x=\frac{2}{x+1}\]

OpenStudy (anonymous):

it loks like this accept -2 is over x

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

Gotcha \[\frac{5}{x^2-1}-\frac{2}{x}=\frac{2}{x+1}\]

OpenStudy (anonymous):

yes thats what it looks like

OpenStudy (anonymous):

So what you want to do firat is factor the denominator of the first fraction so we can figure an LCD

OpenStudy (anonymous):

Ok wish I could help more, but I have to go.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i got x=-3 and x=3 is that right?

OpenStudy (radar):

Just plug your answer back in and check like this